OpenStudy (brainzonly):

Find the general solution for differential equation (D4 - 5D3 + 5D2 + 5D - 6)y = 0 @Kevin Next Q

OpenStudy (brainzonly):

Question closed. Ready?

OpenStudy (kevin):

So we should find D value?

OpenStudy (brainzonly):

Dengan kata lain.... Menemukan solusi umum untuk persamaan diferensial (D4 - 5 d 3 + 5 D 2 + 5 D - 6) y = 0

OpenStudy (brainzonly):

:D My Indonesian is kind of rusty...

OpenStudy (kevin):

Okay I try (D4 - 5D3 + 5D2 + 5D - 6)y = 0 We have (D4 - 5D3 + 5D2 + 5D - 6) = 0 and y = 0 For (D^4 - 5D^3 + 5D^2 + 5D - 6) = 0 (d+1)(d−1)(d−2)(d−3)=0 d+1=0 or d−1=0 or d−2=0 or d−3=0 d=−1 or d=1 or d=2 or d=3

OpenStudy (brainzonly):

:D my solution: Given differential equation, (D4 - 5D3 + 5D2 + 5D - 6)y = 0 => For general solution of equation, Solve D4 - 5D3 + 5D2 + 5D - 6 = 0 => D4 - 5D3 + 6D2 - D2 + 5D - 6 = 0 => D2 (D2 - 5D + 6) - (D2 - 5D + 6) = 0 => (D2 - 5D + 6)(D2 - 1) = 0 ................................(1) Now D2 - 1 = (D - 1)(D + 1) and Factors of D2 - 5D + 6 D2 - 5D + 6 = D2 - 2D - 3D + 6 = D(D - 2) - 3(D - 2) = (D - 3)(D - 2) Therefore, equation (1) implies (D2 - 5D + 6)(D2 - 1) = (D - 3)(D - 2)(D - 1)(D + 1) = 0 => D = 3, 2, 1, -1 or D = -1, 1,, 2, 3 => General solution of differential equation is, => y = C1 e-x + C2 ex + C3 e2x + C4 e3x .

OpenStudy (kevin):

Your Indonesian is good :D

OpenStudy (kevin):

So that's the answer :o

OpenStudy (kevin):

Alright I'm wrong XD Go to another question again. It's very fun

OpenStudy (brainzonly):

Oh YAY! : D